# Multiparty Entanglement Microscopy of Quantum Ising models in 1d, 2d and 3d

### v1

@@ Abstract
Entanglement microscopy reveals the true quantum correlations among the microscopic building blocks of many-body systems [1]. Using this approach, we study the multipartite entanglement of the quantum Ising model in 1d, 2d, and 3d. We first obtain the full reduced density matrix (tomography) of subregions that have at most 4 sites via quantum Monte Carlo, exact diagonalization, and the exact solution in 1d. We then analyze both bipartite and genuine multipartite entanglement (GME) among the sites in the subregion. To do so, we use a variety of measures including the negativity, as well as a true measure of GME: the genuinely multipartite concurrence (or GME concurrence), and its computationally cheaper lower bound, $I_{2}$. We provide a complete proof that $I_{2}$ bounds the GME concurrence, and show how the symmetries of the state simplify its evaluation. For adjacent sites, we find 3- and 4-spin GME present across large portions of the phase diagram, reaching maximum near the quantum critical point. In 1d, we identify the singular scaling of the derivative $dI_{2}/dh$ approaching the critical point. We observe a sharp decrease of GME with increasing dimensionality, coherent with the monogamous nature of entanglement. Furthermore, we find that GME disappears for subregions consisting of non-adjacent sites in both 2d and 3d, offering a stark illustration of the short-ranged nature of entanglement in equilibrium quantum matter [2]. Finally, we analyze the most collective form of entanglement by evaluating the GME concurrence among all spins in the lattice, which can be obtained from a simple observable: the single-site transverse magnetization. This global concurrence is larger in 1d compared to 2d/3d, but it is relatively less robust against perturbations such as local measurements.
@@ II.1 Bipartite entanglement
where $\rho_{A}$ is the RDM of the subregion $A$. Note that we shall use the logarithm base 2 in this paper. In a pure state $|\Psi_{AB}\rangle$, $S^{\rm vN}(A)$ is a valid measure of entanglement between $A$ and its complement, $B$; it fails to be an entanglement measure for mixed states. Although the QMC results are at finite temperature, the computed entropy effectively reflects the groundstate value. This is achieved by scaling the inverse temperature $\beta=1/T$ with the linear system size $L$. $S^{\mathrm{vN}}$ of three adjacent sites are shown in Fig. 2.
@@ II.2 Multipartite entanglement
where the objective function is evaluated by considering two copies of the state $\rho^{\otimes 2}$ on the doubled Hilbert space $(\mathcal{H}_{1}\otimes\cdots\otimes\mathcal{H}_{n})^{\otimes 2}$, on which we maximize over all product states $|\Phi\rangle=|\varphi_{1}\rangle\otimes|\varphi_{2}\rangle\otimes\ldots\otimes|\varphi_{2n}\rangle$. $\Pi$ is the global swap operator that exchanges the two copies, while $\mathcal{P}_{i}=\Pi_{A_{i}}\otimes\mathbbm{1}_{B_{i}}$ is the swap operator for bipartition $\gamma_{i}=(A_{i}|B_{i})$ which exchanges part $A_{i}$ of the two copies. If $I_{2}>0$, there is GME, while no conclusion can be made if $I_{2}\leq 0$. Although $I_{2}$ is weaker than the GMC, it has one important advantage (in addition to lower computational cost): it is obtained by maximization. So even if one finds a non-optimal value that is strictly positive, the state is guaranteed to be GME. In contrast, one cannot rigorously establish that a finite GMC obtained through a numerical convex roof optimization will not be supplanted by a global minimum with $C_{\rm GME}=0$. In practice, we can perform checks on the GMC to gain confidence about the values obtained through the minimization.
@@
where $\rho_{i,j}$ denotes matrix elements of the three-site RDM, and the maximization is performed over local unitary (LU) transformations $\rho\rightarrow U\rho U^{\dagger}$ with $U=U_{1}\otimes U_{2}\otimes U_{3}$. Empirically, we previously employed both $W_{1}$ and a related criterion $W_{2}$ to detect GME in 1d and 2d systems [1]. While both $W_{1}$ and $W_{2}$ successfully detected GME in 1d, they failed to do so in 2d. In contrast, the $I_{2}$ criterion successfully detects GME across 1d, 2d, and 3d, covering both three- and four-spin configurations, see Table 1 for a comparison of $C_{\rm GME}$, $I_{2}$ and $W_{1}$ for three adjacent spins in 1d, 2d and 3d. $I_{2}$ has also been successfully applied to detect GME in other 1d models, such as the XY Ising chain [24] and the cluster-Ising model [38]. These results, combined with its performance across 1d, 2d, and 3d for both ED and QMC states, highlight the robustness of $I_{2}$ as a versatile tool for quantifying GME in complex quantum systems.
@@ IV.2 Decrease of entanglement with separation
Finally, we analyze the decrease of both bipartite and genuine multipartite entanglement as the separation between the subregions $A_{i}$ increases at the QCP. Bipartite entanglement, as measured by negativity, exhibits a rapid decline with separation characterized by the sudden death distance $r_{\mathrm{sd}}$, marking the point where negativity drops to zero. Here the distance is defined as the minimum distance between a site in subregion $A_{1}$ and another in $A_{2}$. Let us begin with the case where each party is made of a single site, which we call the $(1,1)$ configuration, as shown in the inset of Fig. 3 for adjacent sites ($r=1$). Using the exact solution we find that $r^{\text{1d}}_{\text{sd}}=3$ in 1d, in agreement with Ref. [45]. In 2d, the QMC analysis shows that this distance decreases to $r^{\text{2d}}_{\text{sd}}=\sqrt{2}$, while ED on the $5\times 5$ and $6\times 5$ lattices have $r^{\text{2d}}_{\text{sd}}=2$. In 3d, our QMC results yield $r^{\text{3d}}_{\text{sd}}=\sqrt{2}$, consistent with the trend observed in lower dimensions. In the $(2,2)$ configuration, depicted in Fig. 4, the sudden death distance increases to $r^{\text{1d}}_{\text{sd}}=4$ in 1d [45]. In 2d, we consider various separations of two parallel pairs. Using $5\times 5$, $6\times 5$ ED at $h=3$ and $24\times 24$ QMC at $h=3.044$, we get respectively $\mathcal{E}(1)=0.341,0.335,0.344$, $\mathcal{E}(2)=0.00886,0,0.0001$ and $\mathcal{E}(3)=0$ ($6\times 5$ ED), showing a rapid sudden death of the LN (the QMC error bars at $r=2$ include zero). We go further and ask whether the vanishing of the LN also coincides with the vanishing of bipartite entanglement between the two pairs. To do so we evaluate the geometric entanglement
@@
which is the smallest distance with respect to the Frobenius norm, i.e. the Hilbert-Schmidt metric, between the separable set defined by mixtures $\rho_{\rm sep}=\sum_{k}p_{k}\rho_{1}^{k}\otimes\rho_{2}^{k}$, and the density matrix $\rho$; here $\rho^{k}_{i}$ is a valid density matrix for the $i$th pair ($i=1$ or 2). By performing the optimization over the separable states, we find the following upper bounds on $\mathcal{D}$ for the RDM from $6\times 5$ ED: $\mathcal{D}(r\!=\!1)=0.10$, $\mathcal{D}(r\!=\!2)=2.3\times 10^{-4}$, and $\mathcal{D}(r\!=\!3)=2.1\times 10^{-4}$. This suggests that any bound entanglement for $r\geq 2$ would be very weak, if present.
For GME, Table. 2 summarizes $I_{2}$ for various three- and four-spin subregions. Our findings indicate that $I_{2}$ becomes negative as soon as the subregions are disjoint, signifying the absence of GME between disconnected parts. In 1d, GME is detected by $I_{2}$ only for adjacent subregions, and this pattern continues in 2d, where no GME is detected for any disjoint subregion. In 3d, using ED $3\times 3\times 3$, we found no GME in disjoint subregion 126 (see indices in Table 2) and other subregions which involve out-of-plane sites. For the less compact subregions 146, 129, 153, 137, 2468, and 1379, we certify their biseparability using the iterative algorithm introduced in Ref. [49]. A local filtering transformation is applied before the algorithm to maximize the minimum eigenvalue of the RDM. For subregion 126, where $I_{2}$ does not detect GME, we estimated GMC $C_{\rm GME}=1.7\times 10^{-3}$ and $1.1\times 10^{-3}$ for $5\times 5$ and $6\times 5$ ED, respectively. Benchmarking our optimization for the GMC on various states suggests that these values are consistent with a biseparable state.
@@ VI Discussion
Our work invites further investigation along various directions. First, it would be important to investigate larger subregions to understand more collective forms of entanglement. In this respect, it would be desirable to obtain precise RDMs of such subregions with QMC, where a careful handling of the symmetries would facilitate the simulations. Second, it would be natural to employ entanglement microscopy to study the fate of entanglement as a function of temperature. For instance, upper bounds for the temperature at which GME can exist in the 1d Ising model were obtained for 3 adjacent spins [26]. One could extend these results to higher dimensions and to larger subregions. However, this comes with significant computational challenges. Our current practical limit for precision multipartite entanglement microscopy in 3d systems is $L\approx 4$, with a required precision for the RDM matrix elements on the order of $10^{-4}$ to $10^{-5}$. Pushing beyond this limit to larger systems will require more advanced techniques. One promising approach is to integrate tensor-network contraction into Monte Carlo sampling, as demonstrated for entanglement entropy in 2d systems [48]. This method can achieve machine-precision RDM elements for each Monte Carlo configuration. Extending such techniques to 3d systems would enable more precise entanglement studies in higher dimensions and larger subregions, offering deeper insights into the collective behavior of entanglement.

### v2

@@ Abstract
Entanglement microscopy reveals the true quantum correlations among the microscopic building blocks of many-body systems [1]. Using this approach, we study the multipartite entanglement of the quantum Ising model in 1d, 2d, and 3d. We first obtain the full reduced density matrix (tomography) of subregions that have at most 4 sites via quantum Monte Carlo, exact diagonalization, and the exact solution in 1d. We then analyze both bipartite and genuine multipartite entanglement (GME) among the sites in the subregion. To do so, we use a variety of measures including the negativity, as well as a true measure of GME: the genuinely multipartite concurrence (or GME concurrence), and its computationally cheaper lower bound, $I_{2}$. We provide a complete proof that $I_{2}$ bounds the GME concurrence, and show how the symmetries of the state simplify its evaluation. For adjacent sites, we find 3- and 4-spin GME present across large portions of the phase diagram, reaching maximum near the quantum critical point. In 1d, we identify the singular scaling of the derivative $dI_{2}/dh$ approaching the critical point. We observe a sharp decrease of GME with increasing dimensionality, consistent with the monogamous nature of entanglement. Furthermore, we find that GME vanishes for subregions consisting of non-adjacent sites in both 2d and 3d, offering a stark illustration of the short-ranged nature of entanglement in equilibrium quantum matter [2]. Finally, we analyze the most collective form of entanglement by evaluating the GME concurrence among all spins in the lattice, which can be obtained from a simple observable: the single-site transverse magnetization. This global concurrence is larger in 1d compared to 2d/3d, but it is relatively less robust against perturbations such as local measurements.
@@ II.1 Bipartite entanglement
where $\rho_{A}$ is the RDM of the subregion $A$. Note that we shall use the logarithm base 2 in this paper. In a pure state $|\Psi_{AB}\rangle$, $S^{\rm vN}(A)$ is a valid measure of entanglement between $A$ and its complement, $B$; it fails to be an entanglement measure for mixed states. In finite temperature QMC simulations, one usually scales the inverse temperature $\beta=1/T$ with the linear system size $L$ to reduce thermal effects and obtain the groundstate information. $S^{\mathrm{vN}}$ of three adjacent sites are shown in Fig. 2.
@@ II.2 Multipartite entanglement
where the objective function to be maximized is evaluated by considering two copies of the state $\rho^{\otimes 2}$ on the doubled Hilbert space $(\mathcal{H}_{1}\otimes\cdots\otimes\mathcal{H}_{n})^{\otimes 2}$, on which we maximize over all product states $|\Phi\rangle=|\varphi_{1}\rangle\otimes|\varphi_{2}\rangle\otimes\ldots\otimes|\varphi_{2n}\rangle$. $\Pi$ is the global swap operator that exchanges the two copies, while $\mathcal{P}_{i}=\Pi_{A_{i}}\otimes\mathbbm{1}_{B_{i}}$ is the swap operator for bipartition $\gamma_{i}=(A_{i}|B_{i})$ which exchanges part $A_{i}$ of the two copies. If $I_{2}>0$, there is GME, while no conclusion can be made if $I_{2}\leq 0$. Although $I_{2}$ is weaker than the GMC, it has one important advantage (in addition to lower computational cost): it is obtained by maximization. So even if one finds a non-optimal value that is strictly positive, the state is guaranteed to be GME. In contrast, one cannot rigorously establish that a finite GMC obtained through a numerical convex roof optimization will not be supplanted by a global minimum with $C_{\rm GME}=0$. In practice, we can perform checks on the GMC to gain confidence about the values obtained through the minimization.
@@
where $\rho_{i,j}$ denotes matrix elements of the three-site RDM, and the maximization is performed over local unitary (LU) transformations $\rho\rightarrow U\rho U^{\dagger}$ with $U=U_{1}\otimes U_{2}\otimes U_{3}$. The criterion for multipartite entanglement is $W_{1}>0$. Empirically, we previously employed both $W_{1}$ and a related criterion $W_{2}$ [37] to detect GME in 1d and 2d systems [1]. While both $W_{1}$ and $W_{2}$ successfully detected GME in 1d, they failed to do so in 2d. In contrast, the $I_{2}$ criterion successfully detects GME across 1d, 2d, and 3d, covering both three- and four-spin configurations, see Table 1 for a comparison of $C_{\rm GME}$, $I_{2}$ and $W_{1}$ for three adjacent spins in 1d, 2d and 3d. $I_{2}$ has also been successfully applied to detect GME in other 1d models, such as the XY Ising chain [24] and the cluster-Ising model [38]. These results, combined with its performance across 1d, 2d, and 3d for both exact diagonalization (ED) and QMC states, highlight the robustness of $I_{2}$ as a versatile tool for quantifying GME in complex quantum systems.
@@ IV.2 Decrease of entanglement with separation
Finally, we analyze the decrease of both bipartite and genuine multipartite entanglement as the separation between the subregions $A_{i}$ increases at the QCP. Bipartite entanglement, as measured by negativity, exhibits a rapid decline with separation characterized by the sudden death distance $r_{\mathrm{sd}}$, marking the point where negativity drops to zero. Here the distance is defined as the minimum distance between a site in subregion $A_{1}$ and another in $A_{2}$. Let us begin with the case where each party is made of a single site, which we call the $(1,1)$ configuration, as shown in the inset of Fig. 3 for adjacent sites ($r=1$). Using the exact solution we find that $r^{\text{1d}}_{\text{sd}}=3$ in 1d, in agreement with Ref. [45]. In 2d, the QMC analysis shows that this distance decreases to $r^{\text{2d}}_{\text{sd}}=\sqrt{2}$, while ED on the $5\times 5$ and $6\times 5$ lattices have $r^{\text{2d}}_{\text{sd}}=2$. In 3d, our QMC results yield $r^{\text{3d}}_{\text{sd}}=\sqrt{2}$, consistent with the trend observed in lower dimensions. In the $(2,2)$ configuration, depicted in Fig. 4, the sudden death distance increases to $r^{\text{1d}}_{\text{sd}}=4$ in 1d [45]. In 2d, we consider various separations of two parallel pairs. Using $5\times 5$, $6\times 5$ ED at $h=3$ and $24\times 24$ QMC at $h=3.044$, we get respectively $\mathcal{E}(1)=0.341,0.335,0.344$, $\mathcal{E}(2)=0.00886,0,0.0001$ and $\mathcal{E}(3)=0$ ($6\times 5$ ED), showing a rapid sudden death of the LN (the QMC error bars at $r=2$ include zero). We go further and ask whether the vanishing of the LN also coincides with the vanishing of bipartite entanglement between the two pairs. To do so we evaluate the geometric entanglement [49]
@@
which is the smallest distance with respect to the Frobenius norm, i.e. the Hilbert-Schmidt metric, between the separable set defined by mixtures $\rho_{\rm sep}=\sum_{k}p_{k}\rho_{1}^{k}\otimes\rho_{2}^{k}$, and the density matrix $\rho$; here $\rho^{k}_{i}$ is a valid density matrix for the $i$th party ($i=1$ or 2). By performing the optimization over the separable states using the Gilbert algorithm [50, 51], we find the following upper bounds on $\mathcal{D}$ for the RDM from $6\times 5$ ED: $\mathcal{D}(r\!=\!1)=0.10$, $\mathcal{D}(r\!=\!2)=2.3\times 10^{-4}$, and $\mathcal{D}(r\!=\!3)=2.1\times 10^{-4}$ [^1] Ref. [50] choses random product states at each iteration to update the separable state. In contrast, we follow the original deterministic protocol of Gilbert [51], which performs an optimization at each iteration to find a product state, leading to faster convergence.. We found that the resulting approximate closest separable state does not yield an entanglement witness, so these values are not fully converged, and can be further reduced with additional work. Nevertheless, the small distances suggest that any bound entanglement for $r\geq 2$ would be very weak, if present.
For GME, Table. 2 summarizes $I_{2}$ for various three- and four-spin subregions. Our findings indicate that $I_{2}$ becomes negative when the subregions are disjoint, signifying the absence of GME between disconnected parts. In 1d, GME is detected by $I_{2}$ only for adjacent subregions, and this pattern continues in 2d, where no GME is detected for any disjoint subregion. In 3d, using ED $3\times 3\times 3$, we found no GME in disjoint subregion 126 (see indices in Table 2) and other subregions which involve out-of-plane sites. For the less compact subregions 146, 129, 153, 137, 2468, and 1379, we certify their biseparability using the iterative algorithm introduced in Ref. [48]. A local filtering transformation is applied before the algorithm to maximize the minimum eigenvalue of the RDM. For subregion 126, where $I_{2}$ does not detect GME, we estimated GMC $C_{\rm GME}=1.7\times 10^{-3}$ and $1.1\times 10^{-3}$ for $5\times 5$ and $6\times 5$ ED, respectively. Benchmarking our optimization for the GMC on various states suggests that these values are consistent with a biseparable state.
@@ VI Discussion
We began systematically charting the structure of bipartite and multipartite entanglement in the quantum-Ising models in 1d, 2d, and 3d via entanglement microscopy. For 4 sites or less, all forms of entanglement between nearby subregions (parties) peak near the quantum critical point. When increasing the separation between the parties, we find that both two-party and multi-party entanglement are extremely short-ranged. In particular, GME is shared among neighbouring subregions only. This demonstrates that, even at criticality, long-range correlations are predominantly classical. By increasing the spatial dimension, monogamy suppresses entanglement, with an even stronger impact in the multipartite case. We further considered the most collective form of multipartite entanglement involving all spins, and showed that it reduces to the most local quantity—the entanglement between a single site and the rest of the system, which can be readily computed with entanglement microscopy. We have further shown that although the collective form is the largest in the ordered phase, a single measurement can drastically reduce it, and the post-measurement multipartite concurrence peaks near the transition. These findings establish entanglement microscopy as a practical tool for diagnosing phases and criticality, and pave the way for applying the approach to more exotic models where entanglement may organize itself in fundamentally different ways.
Our work invites further investigation along various directions. First, it would be important to investigate larger subregions to understand more collective forms of entanglement. In this respect, it would be desirable to obtain precise RDMs of such subregions with QMC, where a careful handling of the symmetries would facilitate the simulations. Second, it would be natural to employ entanglement microscopy to study the fate of entanglement as a function of temperature. For instance, upper bounds for the temperature at which GME can exist in the 1d Ising model were obtained for 3 adjacent spins [26]. One could extend these results to higher dimensions and to larger subregions. However, this comes with significant computational challenges. Our current practical limit for precision multipartite entanglement microscopy in 3d systems is $L\approx 4$, with a required precision for the RDM matrix elements on the order of $10^{-4}$ to $10^{-5}$. Pushing beyond this limit to larger systems will require more advanced techniques. One promising approach is to integrate tensor-network contraction into Monte Carlo sampling, as demonstrated for entanglement entropy in 2d systems [55]. Extending such techniques to 3d systems would enable more precise entanglement studies in higher dimensions and larger subregions, offering deeper insights into the collective behavior of entanglement.